i dont understand how to find x in
3x/4 - x/3 + 1 = 4x/5 - 3/20

Answer by stanbon(57967)

3x/4 - x/3 + 1 = 4x/5 - 3/20
---
Multiply thru by 60 to get:
45x - 20x + 60 = 48x - 9
----
25x + 60 = 48x - 9
----
23x = 69
x = 3
============
Cheers,
Stan H.
============